# Growing spines ad infinitum

### v1

@@ Abstract.
ADM was partially supported by GNSAGA (INdAM, Italian National Institute of High Mathematics) and PRIN 2022 “Models, sets and classification”.
@@ 1. Introduction
In this paper, we study the model theory of ordered abelian groups, building on the work of Schmitt [21] as well as Cluckers and Halupczok [3]. Ordered abelian groups are much-studied objects in model theory, with many ongoing developments. For example, there are several recent results on the classification of ordered abelian groups according to dividing lines ([14], [2], [7]), elimination of imaginaries [25], and pairs of models (e.g. [4], [9]). The study of ordered abelian groups is often motivated by that of valued fields, since by the Ax-Kochen/Ershov principle, understanding value groups is a key step in order to understand valued fields. This connection is explicit in e.g. [2], [10], and [24]. Properties of value groups are also often used to define henselian valuations in the language of rings, see [12] and [15].
@@
The paper is organized as follows. In Section 2, we give an introduction to spines, as well as the language $\mathcal{L}_{syn}$. We study the behaviour of convex subgroups in this language and show that the spines of a convex subgroup $H\subseteq G$ embed into the spines of $G$ (Corollary 2.7). Spines can be seen as coloured multi-orders, a multisorted generalization of linear orders (see Definition 3.1). Section 3 hence discusses coloured multi-orders. In particular, we show that given an elementary pair $A\preccurlyeq B$ of coloured multi-orders, the convex hull of $A$, as well as the left and right closures of $A$, are elementary substructures of $B$ (Lemma 3.5). In the remainder of Section 3, we record some facts about augmentability of coloured multi-orders for future reference. Most notably, in Lemma 3.8, we characterize augmentability of a coloured multi-order $A$ on the right (resp. left): We show that there is a coloured multi-order $B$ such that $A\preccurlyeq A+B$ (resp. $A\preccurlyeq B+A$) holds if and only if some component of $A$ is unbounded on the right (resp. left). In Section 4, we prove our main result: any non-trivial ordered abelian group is augmentable by infinites (Theorem 4.3), and discuss how to tell that the quotient formed by the infinites is ($p$-)divisible (Corollary 4.5). As an application of our main result, we discuss consequences for henselian valued fields in Section 5, in particular Proposition 5.1 and Theorem 5.5. In the appendix, we give an alternative proof of characterizing which coloured multi-orders are augmentable (Corollary A.3).
@@ Acknowledgements
We thank Sylvy Anscombe, Philip Dittmann, Arno Fehm, Immanuel Halupczok, Mathias Stout, and Floris Vermeulen for their useful comments and discussions.
@@ 2. Spines
Since $\mathcal{A}$ is a closed set of sorts, it is stably embedded and pure in the sense of [19, Appendix A]. This result of quantifier elimination says that any $\mathcal{L}_{\text{syn}}$-formula $\phi(\bar{x},\bar{\eta})$, with $G$-variables $\bar{x}$ and $\mathcal{A}$-variables $\bar{\eta}$, is a boolean combination of formulas of the form
  • $\psi(\bar{x})$, where $\psi$ is quantifier free in the language
    $(G,0,+,-,<,(\equiv_{m})_{m\in\mathbb{N}},(=_{\bullet}k_{\bullet})_{k\in\mathbb{N}},(\equiv_{\bullet m}k_{\bullet})_{m,k\in\mathbb{N}},(D^{[p^{s}]}_{p^{r}})_{p\in\mathcal{P},s\geqslant r\in\mathbb{N}}),$
@@
  • $\chi(\bar{x},\bar{\eta}):=\xi((\mathfrak{s}_{p^{r}}(\sum_{i<n}z_{i}x_{i}),\mathfrak{t}_{p}(\sum_{i<n}z_{i}x_{i}))_{p\in\mathbb{P},r\in\mathbb{N},z_{0},\dots,z_{n-1}\in\mathbb{Z}},\bar{\eta})$, where $\xi$ is an $\mathcal{A}$-formula.
It follows in particular that a sentence in $\mathcal{L}$ must be a sentence in $\mathcal{A}$.
@@
We call this fact the “CH/S transfer principle” as an echo to the AK/E transfer principle, after Schmitt, who stated in very different terms a similar transfer principle in his habilitation thesis, and Cluckers-Halupczok, who derived the language $\mathcal{L}_{\mathrm{syn}}$ for quantifier elimination, giving us this more refined version of the transfer principle.
In general, given $G^{\prime}$ an ordered abelian group and $G\leq G^{\prime}$ a subgroup, we do not have in general an inclusion of spines, that is, it might be that $group\mathcal{A}(G)$ is not a subset of $\mathcal{A}(G^{\prime})$. For example, $\mathbb{Z}\subseteq\mathbb{Q}$, but $\mathcal{S}_{p}(\mathbb{Z})$ has 1 point (for $\{0\}$) where $\mathcal{S}_{p}(\mathbb{Q})$ is empty.
When we restrict our study to convex subgroup, the story changes:
@@
    Clearly $G_{\alpha}\subseteq G_{\alpha^{\prime}}$, since the union is taken over a larger set of indices, and since as we’ve just seen $\mathfrak{s}_{p}$ and $\mathfrak{s}_{p}^{\prime}$ agree on $H$. Now assume, towards a contradiction, that this inclusion is strict. This means that there is $b\in G\setminus pG$ such that $a\notin G_{\mathfrak{s}_{p}^{\prime}(b)}$ and $G_{\mathfrak{t}_{p}(a)}\subsetneq G_{\mathfrak{s}_{p}^{\prime}(b)}$. By Lemma 2.5, we can assume $b\in H$. But then $G_{\mathfrak{s}_{p}^{\prime}(b)}=G_{\mathfrak{s}_{p}(b)}$ appears in the definition of $G_{\mathfrak{t}_{p}(a)}$, that is, $G_{\mathfrak{s}_{p}(b)}\subseteq G_{\mathfrak{t}_{p}(a)}$, a contradiction.
@@
Let $G$ be an ordered abelian group and $H$ be a convex subgroup of $G$. We equip $G$ and $H$ with the language $\mathcal{L}_{\mathrm{syn}}$ independently of one another. Then $H$ can be seen as an $\mathcal{L}_{\mathrm{syn}}$-substructure of $G$. More precisely, we define an embedding $\iota\colon H\rightarrow G$ extending the inclusion $H\subseteq G$ by setting, for any $a\in H$ and any prime $p$, $\iota(\mathfrak{s}_{p}^{H}(a))=\mathfrak{s}_{p}^{G}(a)$, $\iota(\mathfrak{t}_{p}^{H}(a))=\mathfrak{t}_{p}^{G}(a)$ and $\iota({\mathfrak{t}_{p}^{+}}^{H}(a))={\mathfrak{t}_{p}^{+}}^{G}(a)$. In this way, $\iota(H)$ is $H$ equipped with the restriction of relations and symbols of $\mathcal{L}_{\mathrm{syn}}$ from $G$ to $H$.
@@
The fact that $\iota$ is well-defined follows from Lemma 2.6 and Lemma 2.5. Remains to check that $\iota$ is an $\mathcal{L}_{\mathrm{syn}}$-embedding: this is clear, since all predicates represent “local” first order properties of subgroups of the form $G_{\alpha}$, and these are computed in $H$ exactly as in $G$. ∎
We can even be more precise: since points in the spines of $H$ do not change when computed in $G$ – except for potential initial points in $\mathcal{T}_{p}^{+}$ – we have the following:
Corollary 2.8.
Let $G$ be an ordered abelian group and $H$ a convex subgroup of $G$. We equip $G$ and $H$ with a copy of $\mathcal{L}_{\mathrm{syn}}$. Let $\mathcal{C}=\left\{\alpha\in\mathcal{A}(H)\mid G_{\alpha}=H\right\}$. Then $\mathcal{A}(H)\setminus\mathcal{C}$ is an end segment of $\mathcal{A}(G)$.
@@ 3. Coloured multi-orders
The following lemma is an immediate generalization to multi-orders of [20, Theorem 2.1], also obtained using Ehrenfeucht-Fraïssé games:
@@
Let $a\in A$. With the notation above, we have $B_{\leqslant a}=\left\{x\in B\mid x\leqslant a\right\}$ and $A_{>a}=\left\{x\in A\mid x>a\right\}$. Denote $B_{\leqslant a}+A_{>a}$ by $C_{a}$. We have naturally $A_{\leqslant a}\preccurlyeq B_{\leqslant a}$, therefore by Lemma 3.3, $A=A_{\leqslant a}+A_{>a}\preccurlyeq C_{a}\preccurlyeq B$. The chain of structures $(C_{a})_{a\in A}$ is thus elementary, and by Tarski’s chain lemma [22, Theorem 2.1.4 & Exercise 2.1.1], $A\preccurlyeq\bigcup_{a\in A}C_{a}\preccurlyeq B$, and $\bigcup_{a\in A}C_{a}$ is exactly $B_{\leqslant A}$.
@@
Now fix $C^{\prime}\subseteq C$ and let $\lambda=|C^{\prime}|$ enumerate $C^{\prime}$. Fix $A_{0}=A$ and for each $i<\lambda$, let $A_{i+1}$ be an elementary extension of $A_{i}$ realizing only the cut number $i$ as explained above. For limit ordinals $i<\lambda$, let $A_{i}=\bigcup_{j<i}A_{j}$. Now by Tarski’s chain lemma [22, Theorem 2.1.4] $A_{\lambda}=\bigcup_{i<\lambda}A_{i}$ is an elementary extension of $A$, and it realizes exactly the cuts in $C^{\prime}$. ∎
@@ 4. Augmentability by infinites
Since $G\preccurlyeq G^{\prime}$, clearly $\mathcal{A}(G)\preccurlyeq\mathcal{A}(G^{\prime})$. By Corollary 2.7, $H$ is an $\mathcal{L}_{\mathrm{syn}}$-substructure of $G^{\prime}$.
@@
Proof of claim.
Indeed, by Corollary 2.8, $\mathcal{A}(H)\setminus\mathcal{C}$ is an end segment of $\mathcal{A}(G^{\prime})$, and by definition of the right hull, so is $\mathcal{B}$. It is clear that points in $\mathcal{B}$ correspond to convex subgroups of $H$, that is, $\mathcal{B}\subseteq\mathcal{A}(H)\setminus\mathcal{C}$.
@@
Now $(H,\mathcal{C}+\mathcal{B})$, together with the restriction of $\mathfrak{t}_{p},\mathfrak{s}_{p},D^{[p^{s}]}_{p^{r}},=_{\bullet}k_{\bullet},\equiv_{\bullet}k_{\bullet}$ to $H$, is an $\mathcal{L}_{syn}$-substructure of $G^{\prime}$, and similarly $(G,\mathcal{A}(G))$ is an $\mathcal{L}_{syn}$-substructure of $(H,\mathcal{C}+\mathcal{B})$. By Lemma 3.5, $\mathcal{A}(G)\setminus\mathcal{C}\preccurlyeq\mathcal{B}\preccurlyeq\mathcal{A}(G^{\prime})\setminus\mathcal{C}$. Thus by Lemma 3.3
@@
By 2.4, $G\preccurlyeq H\preccurlyeq G^{\prime}$. ∎

### v3

@@ Abstract.
ADM was partially supported by GNSAGA (INdAM, Italian National Institute of High Mathematics) - Project “Model theory and applications to algebra and combinatorics”, and PRIN 2022 “Models, sets and classification”.
@@
0. A note on a former result
This paper was written without knowledge of the paper “Inclusions et produits de groupes abéliens ordonnés étudiés au premier ordre” by Françoise Delon and François Lucas. In this paper from 1989, they prove, among other results, that given $G\preccurlyeq G^{*}$ ordered abelian groups and $H=\langle G\rangle$ the convex hull of $G$, then $G\preccurlyeq H\preccurlyeq G^{*}$, that is, our Lemma 4.1 is exactly Proposition 1 from [5].
In order to prove this, they work in the formalism of Schmitt, see [26]. They argue in Example 1 that given coloured linear orders $C_{1}\preccurlyeq C_{2}$, if $C_{3}$ is either the initial closure of $C_{1}$ in $C_{2}$, the final closure of $C_{1}$ in $C_{2}$, or the convex hull of $C_{1}$ in $C_{2}$, then $C_{1}\preccurlyeq C_{3}\preccurlyeq C_{2}$. This is essentially what we prove in Lemma 3.5, although we are doing it for multi-orders instead of linear orders.
Their Lemma 1-1 is saying that if $H$ is a convex subgroup of an ordered abelian group $G$, then $Sp_{n}(H)$ is an initial segment of $Sp_{n}(G)$, where $Sp_{n}$ denote the Spectrum of order $n$, a structure coming from Schmitt’s formalism, and very similar to the $n$-spines of Cluckers-Halpczok’s language. This result is therefore very similar to our Corollary 2.9.
It is then just a matter of piecing it together to obtain Proposition 1 from their paper or Lemma 4.1 of this paper.
Because we didn’t know that their result existed, we ended up reproving it ourselves independently, going by a very similar argument but using Cluckers-Halupczok’s language instead of Schmitt’s.
Our main theorem, Theorem 4.3, is however new, to the extent of our knowledge. Its application to definability of valuations is also a new result.
@@ 1. Introduction
In this paper, we study the model theory of ordered abelian groups, building on the work of Schmitt [26] as well as Cluckers and Halupczok [3]. Ordered abelian groups are much-studied objects in model theory since the pioneer work of Robinson and Zakon [24]. Among important and classical contributions, we may cite [5] (on elementary subgroups), [27] (quantifiers elimination relative to spines) and [19] (ordered and valued divisible groups). Similar methods were developed in the context of valued vector spaces in [18] and applied in [20]. There are several recent developments on the classification of ordered abelian groups according to dividing lines ([15], [2], [8]), elimination of imaginaries [31], and pairs of models (e.g. [4], [10]). The study of ordered abelian groups is often motivated by that of valued fields, since by the Ax-Kochen/Ershov principle, understanding value groups is a key step in order to understand valued fields. This connection is explicit in e.g. [2], [11], and [30]. Properties of value groups are also often used to define henselian valuations in the language of rings, see [13] and [16].
@@
The paper is organized as follows. In Section 2, we give an introduction to spines, as well as the language $\mathcal{L}_{\text{syn}}$. We study the behaviour of convex subgroups in this language, and discuss when the spines of a convex subgroup $H\subseteq G$ can be embedded into the spines of $G$ (Corollary 2.7). Spines can be seen as coloured multi-orders, a multisorted generalization of linear orders (see Definition 3.1). Section 3 hence discusses coloured multi-orders. In particular, we show that given an elementary pair $A\preccurlyeq B$ of coloured multi-orders, the convex hull of $A$, as well as the left and right closures of $A$, are elementary substructures of $B$ (Lemma 3.5). In the remainder of Section 3, we record some facts about augmentability of coloured multi-orders for future reference. Most notably, in Lemma 3.8, we characterize augmentability of a coloured multi-order $A$ on the right (resp. left): We show that there is a coloured multi-order $B$ such that $A\preccurlyeq A+B$ (resp. $A\preccurlyeq B+A$) holds if and only if some component of $A$ is unbounded on the right (resp. left). In Section 4, we prove our main result: any non-trivial ordered abelian group is augmentable by infinites (Theorem 4.3), and discuss how to tell that the quotient formed by the infinites is ($p$-)divisible (Corollary 4.5). As an application of our main result, we discuss consequences for henselian valued fields in Section 5, in particular Proposition 5.1 and Theorem 5.5. In the appendix, we give an alternative proof of characterizing which coloured multi-orders are augmentable (Corollary A.3).
@@ Acknowledgements
We deeply thank Salma Kuhlmann for pointing out to us the work of Delon and Lucas [5]. We thank Sylvy Anscombe, Philip Dittmann, Arno Fehm, Immanuel Halupczok, Mathias Stout, and Floris Vermeulen for their useful comments and discussions.
@@ 2. Spines
Since $\mathcal{A}$ is a closed set of sorts, it is stably embedded and pure in the sense of [23, Appendix A]. This result of quantifier elimination says that any $\mathcal{L}_{\text{syn}}$-formula $\phi(\bar{x},\bar{\eta})$, with $G$-variables $\bar{x}$ and $\mathcal{A}$-variables $\bar{\eta}$, is a boolean combination of formulas of the form
  • $\psi(\bar{x})$, where $\psi$ is quantifier-free in the language
    $(G,0,+,-,<,(\equiv_{m})_{m\in\mathbb{N}},(=_{\bullet}k_{\bullet})_{k\in\mathbb{N}},(\equiv_{\bullet m}k_{\bullet})_{m,k\in\mathbb{N}},(D^{[p^{s}]}_{p^{r}})_{p\in\mathbb{P},s\geqslant r\in\mathbb{N}}),$
@@
  • $\chi(\bar{x},\bar{\eta}):=\xi((\mathfrak{s}_{p_{j}}(\sum_{i<n}z_{i}^{j}x_{i}),\mathfrak{t}_{p_{j}}(\sum_{i<n}z_{i}^{j}x_{i}))_{z_{0}^{j},\dots,z_{n-1}^{j}\in\mathbb{Z},j\in J},\bar{\eta})$, where $\xi$ is an $\mathcal{L}_{\text{syn}|_{\mathcal{A}}}$-formula, and $J$ is a set of primes (with potential repetitions).
@@
We call this fact the “CH/S transfer principle” as an echo to the AK/E transfer principle, after Schmitt, who stated in very different terms a similar transfer principle in his habilitation thesis [26], and Cluckers-Halupczok, who derived the language $\mathcal{L}_{\mathrm{syn}}$ for quantifier elimination [3], giving us this more refined version of the transfer principle.
In general, given $G^{\prime}$ an ordered abelian group and $G\leq G^{\prime}$ a subgroup, we do not have in general an inclusion of spines, that is, it might be that $\mathcal{A}(G)$ is not a subset of $\mathcal{A}(G^{\prime})$. For example, $\mathbb{Z}\subseteq\mathbb{Q}$, but $\mathcal{S}_{p}(\mathbb{Z})$ has 1 point (for $\{0\}$) where $\mathcal{S}_{p}(\mathbb{Q})$ is empty.
We will now derive some conditions under which (convex) subgroups can be seen as $\mathcal{L}_{\mathrm{syn}}$-substructures.
@@
    Clearly $G_{\alpha}\subseteq G_{\alpha^{\prime}}$, since the union is taken over a larger set of indices, and since by the proof of (1) above $\mathfrak{s}_{p}$ and $\mathfrak{s}_{p}^{\prime}$ agree on $H$. Now assume, towards a contradiction, that this inclusion is strict. This means that there is $b\in G\setminus pG$ such that $a\notin G_{\mathfrak{s}_{p}^{\prime}(b)}$ and $G_{\mathfrak{t}_{p}(a)}\subsetneq G_{\mathfrak{s}_{p}^{\prime}(b)}$. By Lemma 2.5, we can assume $b\in H$. But then $G_{\mathfrak{s}_{p}^{\prime}(b)}=G_{\mathfrak{s}_{p}(b)}$ appears in the definition of $G_{\mathfrak{t}_{p}(a)}$, that is, $G_{\mathfrak{s}_{p}(b)}\subseteq G_{\mathfrak{t}_{p}(a)}$, a contradiction.
@@
Let $G$ be an ordered abelian group and $H$ be a convex subgroup of $G$. We equip $G$ and $H$ with the language $\mathcal{L}_{\mathrm{syn}}$ independently of one another. Assume the following holds:
($\star$) For all primes p and for all aH, if $\text{For all primes $p$ and for all $a\in H$, if }G_{{\mathfrak{t}_{p}^{+}}^{H}(a)}=H\text{ then }G_{{\mathfrak{t}_{p}^{+}}^{G}(a)}=G.$
Equivalently, for all prime $p$ such that there is $a\in H$ with $G_{{\mathfrak{t}_{p}^{+}}^{H}(a)}=H$, $G/H$ is divisible by $p$. Then $H$ is naturally an $\mathcal{L}_{\mathrm{syn}}$-substructure of $G$. More precisely, we define an embedding ${\iota\colon H\rightarrow G}$ extending the inclusion $H\subseteq G$ by setting, for any $a\in H$ and any prime $p$, $\iota(\mathfrak{s}_{p}^{H}(a))=\mathfrak{s}_{p}^{G}(a)$, $\iota(\mathfrak{t}_{p}^{H}(a))=\mathfrak{t}_{p}^{G}(a)$ and $\iota({\mathfrak{t}_{p}^{+}}^{H}(a))={\mathfrak{t}_{p}^{+}}^{G}(a)$. In this way, $\iota(H)$ is the $\mathcal{L}_{\mathrm{syn}}$-structure obtained by the restricting the symbols of $\mathcal{L}_{\mathrm{syn}}$ from $G$ to $H$.
Example 2.8.
The condition ($\star$2.7) cannot be dropped. For example, consider the groups $G\coloneqq\mathbb{Z}_{(2)}\oplus\mathbb{Z}$ and $H\coloneqq\{0\}\oplus\mathbb{Z}$. The spines of $H$ do not embed in the spines of $G$ precisely because $G_{{\mathfrak{t}_{2}^{+}}^{G}((0,1))}=H$ and $G_{{\mathfrak{t}_{3}^{+}}^{G}((0,1))}=G$, while $G_{{\mathfrak{t}_{2}^{+}}^{H}((0,1))}=G_{{\mathfrak{t}_{3}^{+}}^{H}((0,1))}=H$.
Proof of Corollary 2.7.
The fact that $\iota$ is well-defined follows from Lemma 2.6 and Lemma 2.5. What remains to check that $\iota$ is an $\mathcal{L}_{\mathrm{syn}}$-embedding: All computations of proper convex subgroups of $H$ yield the same convex subgroup when computed in $G$ (cf. Lemma 2.6). The only potential obstruction comes when a point in $\mathcal{T}_{p}^{+}$ defines $H$ (which will then necessarily be an initial point in $\mathcal{T}_{p}^{+}$). Then condition ($\star$2.7) ensures that doing the same computation in $G$ results with $G$ (rather than a proper convex subgroup containing $H$). ∎
We can even be more precise: since points in the spines of $H$ do not change when computed in $G$ – with the possible exception of initial points in $\mathcal{T}_{p}^{+}$ – we have the following:
Corollary 2.9 (See [5, Lemma 1-1]).
Let $G$ be an ordered abelian group and $H$ a convex subgroup of $G$. We consider $G$ and $H$ each as $\mathcal{L}_{\mathrm{syn}}$-structures. Let $\mathcal{C}=\left\{\alpha\in\mathcal{A}(H)\mid G_{\alpha}=H\right\}$. Then $\mathcal{A}(H)\setminus\mathcal{C}$ is an end segment of $\mathcal{A}(G)$.
@@
Note that $H$ embeds into $G$ in the restricted language $\mathcal{L}_{\mathrm{syn}}^{-}$ where the sorts $\mathcal{T}_{p}^{+}$ (and corresponding functions) are dropped. Any point in $\mathcal{T}_{p}^{+}$ which is not initial is already in $\mathcal{T}_{p}$. The set $\mathcal{C}$ contains all initial points of $\mathcal{T}_{p}^{+}$, and so $\mathcal{A}(H)\setminus\mathcal{C}$ naturally embeds into $\mathcal{A}(G)$. ∎
@@ 3. Coloured multi-orders
The following lemma is an immediate generalization to multi-orders of [25, Theorem 2.1], also obtained using Ehrenfeucht-Fraïssé games:
@@
Let $a\in A$. With the notation above, we have $B_{\leqslant a}=\left\{x\in B\mid x\leqslant a\right\}$ and $A_{>a}=\left\{x\in A\mid x>a\right\}$. Denote $B_{\leqslant a}+A_{>a}$ by $C_{a}$. We have naturally $A_{\leqslant a}\preccurlyeq B_{\leqslant a}$, therefore by Lemma 3.3, $A=A_{\leqslant a}+A_{>a}\preccurlyeq C_{a}\preccurlyeq B$. The chain of structures $(C_{a})_{a\in A}$ is thus elementary, and by Tarski’s chain lemma [28, Theorem 2.1.4 & Exercise 2.1.1], $A\preccurlyeq\bigcup_{a\in A}C_{a}\preccurlyeq B$, and $\bigcup_{a\in A}C_{a}$ is exactly $B_{\leqslant A}$.
@@
Now fix $C^{\prime}\subseteq C$ and let $\lambda=|C^{\prime}|$ enumerate $C^{\prime}$. Fix $A_{0}=A$ and for each $i<\lambda$, let $A_{i+1}$ be an elementary extension of $A_{i}$ realizing only the cut number $i$ as explained above. For limit ordinals $i<\lambda$, let $A_{i}=\bigcup_{j<i}A_{j}$. Now by Tarski’s chain lemma [28, Theorem 2.1.4] $A_{\lambda}=\bigcup_{i<\lambda}A_{i}$ is an elementary extension of $A$, and it realizes exactly the cuts in $C^{\prime}$. ∎
@@ 4. Augmentability by infinites
Since $G\preccurlyeq G^{\prime}$, clearly $\mathcal{A}(G)\preccurlyeq\mathcal{A}(G^{\prime})$.
Claim 1.
The pair $G^{\prime}$ and $H$ satisfies ($\star$2.7) from Corollary 2.7.
Proof of Claim 1.
Assume $G_{{\mathfrak{t}_{p}^{+}}^{H}(a)}=H$ for some prime $p$ and some $a\in H$. Since $H$ is the convex hull of $G$, there exists $b\in G$ such that $|b|>|a|$.
But now $G_{{\mathfrak{t}_{p}^{+}}^{G}(b)}=G$: indeed, $G_{{\mathfrak{t}_{p}^{+}}^{H}(b)}=H$, then by Lemma 2.6, $G_{{\mathfrak{t}_{p}^{+}}^{G^{\prime}}(b)}\supseteq H$, and since $G\subseteq H$, we have $G_{{\mathfrak{t}_{p}^{+}}^{G^{\prime}}(b)}\supseteq G$. By elementarity, this can only happen if $G_{{\mathfrak{t}_{p}^{+}}^{G}(b)}=G$.
Hence $G_{{\mathfrak{t}_{p}^{+}}^{G^{\prime}}(b)}=G^{\prime}$, using elementarity once again.
Now $a\in G_{{\mathfrak{t}_{p}^{+}}^{G^{\prime}}(b)}$ and $b\in H=G_{{\mathfrak{t}_{p}^{+}}^{H}(a)}\subseteq G_{{\mathfrak{t}_{p}^{+}}^{G^{\prime}}(a)}$, thus we must have ${\mathfrak{t}_{p}^{+}}^{G^{\prime}}(b)={\mathfrak{t}_{p}^{+}}^{G^{\prime}}(a)$ and finally $G_{{\mathfrak{t}_{p}^{+}}^{G^{\prime}}(a)}=G^{\prime}$, proving that the pair $H$ and $G^{\prime}$ satisfies condition ($\star$2.7). ∎
Therefore, by Corollary 2.7, $H$ is an $\mathcal{L}_{\mathrm{syn}}$-substructure of $G^{\prime}$.
@@
Proof of Claim 2.
By Corollary 2.9, $\mathcal{A}(H)\setminus\mathcal{C}$ is an end segment of $\mathcal{A}(G^{\prime})$, and by definition of the right hull, so is $\mathcal{B}$. It is clear that points in $\mathcal{B}$ correspond to convex subgroups of $H$, that is, $\mathcal{B}\subseteq\mathcal{A}(H)\setminus\mathcal{C}$.
@@
Now $(H,\mathcal{C}+\mathcal{B})$, together with the restriction of $\mathfrak{t}_{p},\mathfrak{s}_{p},D^{[p^{s}]}_{p^{r}},=_{\bullet}k_{\bullet},\equiv_{\bullet}k_{\bullet}$ to $H$, is an $\mathcal{L}_{syn}$-substructure of $G^{\prime}$, and similarly $(G,\mathcal{A}(G))$ is an $\mathcal{L}_{syn}$-substructure of $(H,\mathcal{C}+\mathcal{B})$. By Lemma 3.5, we have $\mathcal{A}(G)\setminus\mathcal{C}\preccurlyeq\mathcal{B}\preccurlyeq\mathcal{A}(G^{\prime})\setminus\mathcal{C}$. Thus, by Lemma 3.3, we get
@@
By 2.4, we conclude $G\preccurlyeq H\preccurlyeq G^{\prime}$. ∎